OpenStudy (anonymous):

the fourth term of a geometric sequence is 48 and the seventh term of the sequence is 384. determine the first term of the sequence.

OpenStudy (hoblos):

7th = 4th * r^3 r^3 = 384/48 = 8 r=2 4th=1st* r^3 1st = 48/8 = 6

OpenStudy (anonymous):

Yeah, Hoblos is correct.